# Solving Distributed Flexible Job Shop Scheduling Problems in the Wool Textile Industry with Quantum Annealing

### v1

@@ 5.3 Numerical Results
In order to provide the Contributions 2 and 3 formulated in Section 1, the results obtained from the experiment are analyzed from both a qualitative and a quantitative perspective. The first answers the question, how good the returned solution is in terms of system energy and makespan. The second compares QPU access time and CPU time of SA to assess whether there is a speed advantage in using the quantum annealer to solve the DFJSP with distributed operations. The results are influenced by the hardware characteristics, on different hardware other results might be obtained.
@@
System energy and makespan. The goal of the experiment is to find consistent solutions in the shortest time possible. When sampling from QA or SA, the returned sample set contains 1000 samples, in accordance with the defined number of reads. The sample with the lowest energy in the sample set is selected as the best sample result for the respective problem instance. This is not always the solution corresponding to the global minimum represented by the lowest energy possible for the QUBO. For this use case, however, finding the global minimum is not crucial. A solution is satisfactory if it is consistent and represents a valid schedule plan obtained in a reasonable time. Further optimization steps can then be implemented to minimize the energy of the system and the makespan. Fig. 5 shows a comparison of best solution energies found with QA and SA for instances of different sizes.
@@
For the smallest problem instance, both QA and SA found a low energy of 2 and returned valid solutions. For instances of size 100 and 150, both methods delivered consistent solutions, but SA performed better in terms of energy. For problem instances of size 200 and 250, the best solution returned by quantum annealer had 1 and 2 broken constraints, respectively. This explains the increasing energy values for these problems. This observation is related to the percentage of broken chains that increase with problem size and deteriorate solution quality with the growing number of variables in the problem. These results are consistent with results found in other papers. Carugno et al. [4] found that for JSSP instances with an unknown optimal solution, QA is not able to find the best solution that corresponds to the global minimum.
@@
To calculate the makespan in the makespan function (Eq. 10), the minimum ending time and the maximum ending time of all operations are used. These terms represent the minimum and the maximum makespans possible and are the lower and the upper bound of the makespan range, respectively. In terms of makespan, QA returns solutions in the lower half of the possible makespan range for all problem sizes, except for size 250 that is in the upper half of possible solutions (Fig. 6).
@@
To summarize, QA is able to find consistent solutions for problem instances up to 150 variables. For the instance of size 200, the makespan found is still in the first half of solutions, but 1 constraint is violated. For size 250, the makespan is slightly below the maximum possible value and 2 constraints are violated. Current QA technology is subject to hardware and software limitations, but with the growing number of qubits and increasing connectivity between them, problems with a higher number of variables can be potentially embedded into future QPUs. Additionally, improving QPU stability promises to deliver solutions of better quality and with lower energy of the system.
@@
Calculation time. To determine the QPU access time, the instances from size 50 to 250 are solved with the quantum annealer. Based on these results, a trendline for instances of size 300 up to 400 is derived. In this case, since the rate of change in the data decreases and almost levels out, a logarithmic trendline describes the data the best. The R-square value of 0.9767 is a strong indicator that the logarithmic trendline corresponds to the data. With SA it was possible to solve problem instances of all sizes up to the problem instance of 400 variables. Fig. 7 shows the results that were obtained by both methods.
Although SA performs better for small problem sizes, the CPU calculation time of SA increases exponentially with the increasing problem size. On the other hand, due to hardware limitations, the AdvantageSystem4.1 quantum annealer is capable to embed problems up to a maximum size of 250 variables. However, its advantage is that the calculation time increases at a decreasing rate with problem size. For business problems with a very large number of variables, specific to the DFJSP, a calculation time that changes logarithmically with problem size is a big benefit, compared to the exponential growth of the CPU time.
@@ 5.4 Discussion
In this paper, a step-by-step practical implementation of the DFJSP with distributed operations and its evaluation using D-Wave quantum annealer, based on the real use case from the wool textile industry is presented. First, the problem is described mathematically as a constraint satisfaction problem. Then, the process of reformulating the mathematical problem in QUBO format is discussed in detail, as well as its embedding into the D-Wave quantum annealer QPU and the configuration of QPU solver parameters. Finally, solutions are sampled from the QPU to find low-energy states of the cost function. While formulating the QUBO, particular attention is dedicated to the calculation of the appropriate Lagrange parameters (Subsection 4.4), that increase the weights of the penalty terms. Another aspect that is carefully investigated while configuring the experiment is the chain strength parameter (Subsection 5.2). Both aspects make up Contribution 2.
@@
Another aim of the paper was to investigate whether QA is able to provide a speed advantage for the DFJSP with distributed operations specific to the wool textile industry. To answer this question, both qualitative and quantitative analyses are conducted. Current QA technology allows more qubit usage than classical gate model quantum computers. However, there are still limitations imposed by hard- and software imperfections. With this in mind, the test challenged up to the limit the ability of QA to deliver consistent solutions for large problem instances in the shortest time possible. To conduct the experiment, problem instances from size 50 to 250 are solved in steps of 50 variables with QA and with SA. Additionally, problems up to 400 variables are solved only with SA to increase the size range and determine the calculation time trend. The missing data for QA is extrapolated as a logarithmic trendline.
@@
The qualitative analysis of the solutions returned indicates that for problem instances up to a size of 150 variables, both QA and SA, with technical characteristics as used in this research, deliver consistent solutions, and thus valid production schedules. However SA performs slightly better in terms of returned energy. With an increasing number of variables, hardware imperfections become visible, which manifest themselves in an increased number of broken chains in the returned solutions, and consequently, in solutions with higher energy due to broken constraints. For the problem of size 200, the best solution returned by QA has 1 broken constraint, for the instance of 250 variables 2 constraints were violated. Due to a higher usage of physical qubits needed compared to the number of variables in the original QUBO, the maximum problem instance that could be embedded onto the QPU was of 250 variables. For the problem of size 300 no embedding could be found. In terms of makespan, QA finds solutions for almost problem sizes in the lower half of the possible makespan range, except for size 250 which is the upper half of possible solutions. The makespan determined by SA is comparable, although for a few problems instances SA is able to find solutions that are closer to the global minimum. All solutions calculated by SA are consistent, even for the largest instances in the test.
@@
To tackle the Contribution 3, QPU access time and SA CPU time necessary to calculate the solutions to the problems were measured. The results reveal that QA needs more time than SA for smaller problem instances, but the time needed increases at a decreasing rate with the growing problem size. This is not the case for SA, whose CPU time grows exponentially with the increasing number of variables. The CPU timeline and the trendline of the CPU access time are slightly above the value of 300 variables. If this trend will persist with the new QA hardware generation that is to come, every problem with a size larger than this value would potentially be solved faster than with SA. Thus, the question of the speed advantage can be answered positively: QA can potentially result in a speed advantage for the DFJSP with distributed operations.
@@
Even if the QA technology currently still faces limitations, the new generation of quantum annealers with improved hardware promises to speed up solving complex combinatorial optimization problems like the DFJSP. The measurements show that, given the same experimental parameters and improved hardware, QA could provide a speed-up advantage over SA starting with problem instances slightly larger than 300 variables. To test this assumption, new tests must be performed for larger problem instances, for which only the trend could be extrapolated at the moment, as soon as the new generation of quantum annealers becomes available. More qubits and connections between them, as well as more stable systems, will allow embedding larger problems and potentially finding energies closer to the global minimum. Additionally, research on tuning and configuration of QA parameters is another necessity that will allow to minimize the energy and find the best solution. Determining the strength of the qubit chains is a laborious task that is not only model-dependent, but also instance-dependent, and has a large impact on the quality of the obtained solutions.
@@ 6 Conclusion
Many modern manufacturing enterprises evolved from a single to a multi-factory production environment, as is the case with big production companies in the wool textile industry. Planning systems changed from classical JSSP to DFJSP, with companies facing the challenge of increased frequency and complexity of production planning. Classical heuristics deliver valid results on small problem instances but are unable to find solutions for problem sizes relevant to practice. Therefore, methods are required that yield high quality outcomes while reasonable computation times are ensured. In this paper, the concept of QA-based optimization for the DFJSP with geographically distributed operations is investigated to answer the question whether QA is able to provide a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
To do this, we show a step-by-step implementation of the workflow for solving the DFJSP with QA (Contribution 1). A special attention is devoted to topics that do not have a straightforward approach but a great impact on the solution quality, such as calculating the Lagrange parameters or determining the chain strength. The results show that up to a certain problem size, QA delivers consistent solutions that are not the global minimum solution in the solution space but yield a makespan in the first half of the allowed valid solutions (Contribution 2). For all instances, SA delivers comparable or slightly better results, which are consistent with other research. In terms of calculation time, QA needs more QPU time to find solutions for smaller problem instances in the test, but QPU access time increases at a declining rate with the growing problem size. In contrast, the SA CPU time develops exponentially for problems with a larger number of variables. The results of the experiment show that starting with a certain problem size, QA could potentially offer a speed-up advantage, given that the expected improvements in hard- and software will take place (Contribution 3).
One of the future tasks is to solve larger problems on the next generation of quantum annealers to confirm the results obtained in this experiment. Other topics open for future research are QPU parameter tuning and configuration. There is also a need for additional research on embedding techniques that could minimize the length of the chains and allow the embedding of more QUBO variables into physical qubits, thus solving problems of greater size.

### v3

@@ 5.3 Numerical Results
To provide the Contributions 2 and 3 formulated in Section 1, the results obtained from the experiment are analyzed from both a qualitative and a quantitative perspective. The first answers the question of how good the returned solution is in terms of system energy and makespan. The second compares QPU access time and CPU time of SA to assess whether there is a speed advantage in using the quantum annealer to solve the DFJSP with distributed operations. The hardware characteristics influence the results; different hardware may yield different results.
@@
System energy and makespan. The goal of the experiment is to find consistent solutions in the shortest time possible. When sampling from QA or SA, the returned sample set contains 1000 samples, in accordance with the defined number of reads. The sample with the lowest energy in the sample set is selected as the best sample result for the respective problem instance. This is not always the solution corresponding to the global minimum, which is represented by the lowest energy possible for the QUBO. For this use case, however, finding the global minimum is not crucial. A solution is satisfactory if it is consistent and represents a valid schedule plan that can be obtained in a reasonable amount of time. Further optimization steps can then be implemented to minimize the energy of the system and the makespan. Fig. 5 shows a comparison of best solution energies found with QA and SA for instances of different sizes.
@@
For the smallest problem instance, both QA and SA found a low energy of 2 and returned valid solutions. For instances of size 100 and 150, both methods delivered consistent solutions, but SA performed better in terms of energy. For problem instances of size 200 and 250, the best solution returned by the quantum annealer had 1 and 2 broken constraints, respectively. This explains the increasing energy values for these problems. This observation is related to the percentage of broken chains, which increases with problem size and deteriorates solution quality as the number of variables in the problem grows. These results are consistent with results found in other papers. Carugno et al. [5] found that for JSSP instances with an unknown optimal solution, QA is not able to find the best solution that corresponds to the global minimum.
@@
To calculate the makespan in the makespan function (Eq. 10), the minimum ending time and the maximum ending time of all operations are used. These terms represent the minimum and the maximum makespans possible and are the lower and the upper bound of the makespan range, respectively. In terms of makespan, QA returns solutions in the lower half of the possible makespan range for all problem sizes, except for size 250, which is in the upper half of possible solutions (Fig. 6).
@@
To summarize, QA can find consistent solutions for problem instances up to 150 variables. For the instance of size 200, the makespan found is still in the first half of solutions, but 1 constraint is violated. For size 250, the makespan is slightly below the maximum possible value, and 2 constraints are violated. Current QA technology is subject to hardware and software limitations; however, with the growing number of qubits and increasing connectivity between them, problems with a higher number of variables can potentially be embedded into future QPUs. Additionally, improving QPU stability promises to deliver solutions of better quality and with lower energy of the system.
@@
Calculation time. To determine the QPU access time, the instances from size 50 to 250 are solved with the quantum annealer. Based on these results, a trendline is derived for instances of size 300 to 400. In this case, since the rate of change in the data decreases and levels off, a logarithmic trendline best describes the data. The R-squared value of 0.9767 is a strong indicator that the logarithmic trendline accurately represents the data. With SA, it was possible to solve problem instances of all sizes, up to the 400-variable problem instance. Fig. 7 shows the results that were obtained by both methods.
Although SA performs better for small problem sizes, the CPU calculation time of SA increases exponentially with the increasing problem size. On the other hand, due to hardware limitations, the Advantage System 4.1 quantum annealer is capable of embedding problems up to a maximum size of 250 variables. However, its advantage is that the calculation time increases at a decreasing rate with problem size. For business problems with a very large number of variables, specific to the DFJSP, a calculation time that scales logarithmically with problem size is a significant benefit, compared to the exponential growth of CPU time.
@@ 5.4 Discussion
In this paper, a step-by-step practical implementation of the DFJSP with distributed operations and its evaluation using D-Wave quantum annealer, based on a real use case from the wool textile industry, is presented. First, the problem is described mathematically as a constraint satisfaction problem. Then, the process of reformulating the mathematical problem in QUBO format is discussed in detail, as well as its embedding into the D-Wave quantum annealer QPU and the configuration of QPU solver parameters. Finally, solutions are sampled from the QPU to find low-energy states of the cost function. While formulating the QUBO, particular attention is dedicated to the calculation of the appropriate Lagrange parameters (Subsection 4.4), which increase the weights of the penalty terms. Another aspect that is carefully investigated while configuring the experiment is the chain strength parameter (Subsection 5.2). Both aspects make up Contribution 2.
@@
Another aim of the paper was to investigate whether QA can provide a speed advantage for the DFJSP with operations specific to the wool textile industry. To answer this question, both qualitative and quantitative analyses are conducted. Current QA technology enables the use of more qubits than classical gate model quantum computers. However, there are still limitations imposed by hardware and software imperfections. With this in mind, the test challenged to the limit the ability of QA to deliver consistent solutions for large problem instances in the shortest time possible. To conduct the experiment, problem instances with sizes ranging from 50 to 250 are solved in increments of 50 variables using QA and SA. Additionally, problems with up to 400 variables are solved only with SA to increase the size range and determine the calculation time trend. The missing data for QA is extrapolated as a logarithmic trendline.
@@
The qualitative analysis of the solutions returned indicates that for problem instances with up to 150 variables, both QA and SA, with technical characteristics as used in this research, deliver consistent solutions, and thus valid production schedules. However, SA performs slightly better in terms of returned energy. With an increasing number of variables, hardware imperfections become more visible, manifesting as a higher number of broken chains in the returned solutions, and consequently, in solutions with higher energy due to broken constraints. For the problem of size 200, the best solution returned by QA has 1 broken constraint; for the instance of 250 variables, 2 constraints were violated. Due to a higher usage of physical qubits needed compared to the number of variables in the original QUBO, the maximum problem instance that could be embedded onto the QPU was of 250 variables. For the problem of size 300, no embedding could be found. In terms of makespan, QA finds solutions for almost all problem sizes in the lower half of the possible makespan range, except for size 250, which is in the upper half of the possible solutions. The makespan determined by SA is comparable, although for a few problem instances SA can find solutions that are closer to the global minimum. All solutions calculated by SA are consistent, even for the largest instances in the test.
@@
To tackle the Contribution 3, QPU access time and SA CPU time necessary to calculate the solutions to the problems were measured. The results reveal that QA requires more time than SA for smaller problem instances, but the time needed increases at a decreasing rate as the problem size grows. This is not the case for SA, whose CPU time grows exponentially with the increasing number of variables. The CPU timeline and the trendline of the CPU access time are slightly above the value of 300 variables. If this trend persists with the new QA hardware generation that is to come, every problem with a size larger than this value could be solved faster than with SA. Thus, the question of the speed advantage can be answered positively: QA can potentially result in a speed advantage for the DFJSP with distributed operations.
@@
Even if the QA technology currently faces limitations, the new generation of quantum annealers with improved hardware promises to accelerate the solution of complex combinatorial optimization problems, such as the DFJSP. The measurements show that, given the same experimental parameters and improved hardware, QA can provide a speed-up advantage over SA starting with problem instances that are slightly larger than 300 variables. To test this assumption, new tests must be performed for larger problem instances, for which only the trend could be extrapolated at the moment, as soon as the latest generation of quantum annealers becomes available. More qubits and connections between them, as well as more stable systems, will enable the embedding of larger problems and the discovery of energies closer to the global minimum. Additionally, research on tuning and configuration of QA parameters is another necessity that will enable us to minimize energy and find the optimal solution. Determining the strength of the qubit chains is a laborious task that is not only model-dependent but also instance-dependent and has a significant impact on the quality of the obtained solutions.
@@ 6 Conclusion
Many modern manufacturing enterprises have evolved from a single to a multi-factory production environment, as is the case with big production companies in the wool textile industry. Planning systems changed from classical JSSP to DFJSP, with companies facing the challenge of increased frequency and complexity of production planning. Classical heuristics deliver valid results on small problem instances but are unable to find solutions for problem sizes relevant to practice. Therefore, methods are required that yield high-quality outcomes while ensuring reasonable computation times. In this paper, the concept of QA-based optimization for the DFJSP with geographically distributed operations is investigated to answer the question of whether QA can provide a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
To do this, we show a step-by-step implementation of the workflow for solving the DFJSP with QA (Contribution 1). Special attention is devoted to topics that do not have a straightforward approach but a significant impact on the solution quality, such as calculating the Lagrange parameters or determining the chain strength. The results show that up to a certain problem size, QA delivers consistent solutions that are not the global minimum solution in the solution space but yield a makespan in the first half of the allowed valid solutions (Contribution 2). For all instances, SA delivers comparable or slightly better results, which are consistent with other research. In terms of calculation time, QA needs more QPU time to find solutions for smaller problem instances in the test, but QPU access time increases at a declining rate with the growing problem size. In contrast, the SA CPU time develops exponentially for problems with a larger number of variables. The results of the experiment show that, starting with a certain problem size, QA could potentially offer a speed-up advantage, given that expected improvements in hardware and software are anticipated (Contribution 3).
One of the future tasks is to solve larger problems on the next generation of quantum annealers to confirm the results obtained in this experiment. Other topics open for future research are QPU parameter tuning and configuration. There is also a need for additional research on embedding techniques that can minimize the length of the chains and enable the embedding of more QUBO variables into physical qubits, thereby solving problems of greater size.